# Understanding the Cluster LP for Correlation Clustering

### v2

@@ 4.1 Preclustering
Given a Correlation Clustering instance $(V,E^{+}\uplus E^{-})$, a preclustered instance is defined by a pair $({\mathcal{K}},E_{\mathrm{adm}})$, where ${\mathcal{K}}$ is a partition of $V$ (which can also be viewed as a clustering), and $E_{\mathrm{adm}}\subseteq\binom{V}{2}$ is a set of pairs such that for every $uv\in E_{\mathrm{adm}}$, $u$ and $v$ are not in a same set in ${\mathcal{K}}$.
@@
There are two minor differences between our definition and the one in [CLLN23]. First, we require that ${\mathcal{K}}$ forms a partition; this can be guaranteed by adding singletons. Second, we do not require an edge between two different non-singleton atoms to be non-admissible. Our construction can guarantee this condition, but it is not essential.
@@
The following theorem with a worse bound on $|E_{\mathrm{adm}}|$ was proved in [CLLN23]. We give a cleaner proof of the theorem in Section A; as a byproduct, it achieves a better bound on $|E_{\mathrm{adm}}|$.
@@
We can assume in the preclustered instance $({\mathcal{K}},E_{\mathrm{adm}})$, the edges between two different atoms $K$ and $K^{\prime}$ are all admissible, or all non-admissible. If one edge between them is non-admissible, we can change all other edges to non-admissible edges. This will not change the set of good clusterings, and it will decrease $|E_{\mathrm{adm}}|$.
@@
We apply Theorem 14 to obtain a preclustered instance $({\mathcal{K}},E_{\mathrm{adm}})$, with the unknown good clustering ${\mathcal{C}}^{*}_{1}$. We define $K_{u}$ to be the atom that contains $u$, and $k_{u}=|K_{u}|$. We shall use $N_{\mathrm{adm}}(u)$ to be the set of vertices $v$ such that $uv\in E_{\mathrm{adm}}$; so $N_{\mathrm{adm}}(u)=N_{\mathrm{adm}}(v)$ if $v\in K_{u}$. We further process the good clustering ${\mathcal{C}}^{*}_{1}$ using the following procedure in [CLLN23]. This procedure is not a part of our algorithm; it is only for analysis purpose.
@@ 4.2 Bounded Sub-Cluster LP Relaxation for Preclustered Instances
Following [CLLN23], we form an LP relaxation aiming at finding the good clustering ${\mathcal{C}}^{*}_{1}$. In the LP, we have a variable $y^{s}_{S}$, for every $s\in[n]$, and $S\subseteq V$ of size at most $r$ (recall that $r=\Theta(1/\varepsilon^{12})$), that denotes the number of clusters in ${\mathcal{C}}^{*}_{1}$ of size $s$ containing $S$ as a subset. When $S\neq\emptyset$, there is at most one such cluster and thus $y^{s}_{S}\in\{0,1\}$ indicates if $S$ is a subset of a cluster of size $s$ in ${\mathcal{C}}^{*}_{1}$. For every $S\subseteq V$ of size at most $r$, let $y_{S}:=\sum_{s}y^{s}_{S}$ denote the number of clusters (of any size) in ${\mathcal{C}}^{*}_{1}$ containing $S$ as a subset. Again, if $S\neq\emptyset$, then $y_{S}\in\{0,1\}$ indicates if $S$ is a subset of a cluster in ${\mathcal{C}}^{*}_{1}$. For every $uv\in{V\choose 2}$, we have a variable $x_{uv}$ indicating if $u$ and $v$ are separated or not in ${\mathcal{C}}^{*}_{1}$. We call the LP the bounded sub-cluster LP relaxation, as we have variables indicating if a small set $S$ is a subset of a cluster or not.
@@
(6) gives the definition of $y_{S}$, (7) requires $u$ to be contained in some cluster, and (8) gives the definition of $x_{uv}$. (9) says if $y^{s}_{S}=1$, then there are exactly $s$ elements $u\in V$ with $y^{s}_{Su}=1$. (An exception is when $S=\emptyset$; but the equality also holds.) (10) is the non-negativity constraint. (11) and (12) follows from that ${\mathcal{C}}^{*}_{1}$ is a good clustering, and (13) follows from (A1). The left side of (14) is the number of clusters of size $s$ containing $S$ but does not contain any vertex in $T$. So the inequality holds. This corresponds to a Sherali-Adams relaxation needed for the correlated rounding [RT12], see Lemma 16. The running time for solving the LP is $n^{O(r)}=n^{O(1/\varepsilon^{12})}$.
@@ 4.3 Sampling One Cluster Using LP Solution to the Bounded Sub-Cluster LP
We solve the bounded sub-cluster LP to obtain the $y$ and $x$ vectors. Given $y$, we can use the procedure construct-cluster described in Algorithm 3, which is from [CLLN23], to produce a random cluster $C$.
Algorithm 3 construct-cluster$(y)$
@@
2:randomly choose a vertex $u\in V$, so that $u$ is chosen with probability $\frac{y^{s}_{u}}{sy^{s}_{\emptyset}}$
3:define a vector $y^{\prime}$ such that $y^{\prime}_{S}=\frac{y^{s}_{Su}}{y^{s}_{u}}$ for every $S\subseteq V$ of size at most $r-1$
@@
As in [CLLN23], we define $\mathrm{err}^{s}_{vw|u}$ to be the error generated by the procedure when we choose $s$ as the cardinality and $u$ as the pivot:
$\mathrm{err}^{s}_{vw|u}:=\left|\Pr\big{[}v,w\in C|s,u\big{]}-\frac{y^{s}_{uvw}}{y^{s}_{u}}\right|,\forall vw\in{V\choose 2},$
@@
The following two lemmas can be proved using the same arguments as in [CLLN23].
Lemma 17 ([CLLN23]).
@@
Lemma 18 ([CLLN23]).
@@
    $\Pr\left[|\{v,w\}\subseteq C\right]\leq\frac{1}{y_{\emptyset}}\cdot y_{vw}+\mathrm{err}_{vw}$.
A similar lemma to the following is proved in [CLLN23]. The parameters we use here are slightly different and we provide a proof for completeness.
@@
Throughout the proof, we assume $u,v,w$ are all in $V$, $vw$ and $uw$ are in ($V\choose 2$.
Fix some $s\in[n],u\in V$ with $y^{s}_{u}>0$, and we now bound $\sum_{vw}\mathrm{err}^{s}_{vw|u}$. If $s=k_{u}$, then $C=K_{u}$; no errors will be created and the quantity is 00. Assume $s>k_{u}$. By (13), we have that $s>k_{u}+\varepsilon_{1}\cdot|N_{\mathrm{adm}}(u)|$, since otherwise we shall $y^{s}_{u}=0$. By the second property of Lemma 16, we have $\sum_{vw}\mathrm{err}^{s}_{vw|u}\leq\frac{\varepsilon_{\mathrm{rt}}}{2}|N_{\mathrm{adm}}(u)|^{2}$. (Notice that if one of $v$ and $w$ is not in $N_{\mathrm{adm}}(u)$, then $\mathrm{err}^{s}_{vw|u}=0$.) Recall that $\varepsilon_{\mathrm{rt}}=\varepsilon_{1}^{2}$. Therefore,
@@
$=\frac{\varepsilon_{1}}{2}\cdot|N_{\mathrm{adm}}(u)|\cdot\sum_{v\in N_{\mathrm{adm}}(u)}\frac{y^{s}_{uv}}{y^{s}_{u}}=\frac{\varepsilon_{1}}{2}\cdot\sum_{v,w\in N_{\mathrm{adm}}(u)}\frac{y^{s}_{uv}}{y^{s}_{u}}.$
The first equality is by (9) and $y^{s}_{uv}=y^{s}_{u}$ for every $v\in K_{u}$. (To see this, notice that $y^{s}_{uv}\leq y^{s}_{u}$ is implied by (14). We have $y_{uv}=\sum_{s}y^{s}_{uv}$, $y_{u}=\sum_{s}y^{s}_{u}$, and $y_{uv}=y_{u}=1$ if $v\in K_{u}$.)
@@
$=\varepsilon_{1}\sum_{uw\in E_{\mathrm{adm}}}\Pr[C\cap\{u,w\}\neq\emptyset\mid s].$
To see the last inequality, notice that $\frac{y^{s}_{uv}}{y^{s}_{v}}=\Pr[u\in C|s,v\text{ is pivot}]\leq\Pr[C\cap\{u,w\}\neq\emptyset|s,v\text{ is pivot}]$. The same inequality holds for $\frac{y^{s}_{vw}}{y^{s}_{v}}$.
@@
To see the last inequality, we notice that $C\cap\{u,w\}\neq\emptyset$ is the union of the 3 disjoint events: $u\in C$ and $w\notin C$, $u\notin C$ and $w\in C$, and $\{u,w\}\notin C$. By Lemma 18, we have $\Pr[C\cap\{u,w\}\neq\emptyset]\leq\frac{2x_{vw}+y_{vw}}{y_{\emptyset}}+3\cdot\mathrm{err}_{uw}\leq\frac{2}{y_{\emptyset}}+3\cdot\mathrm{err}_{uw}$. So, we have $\sum_{vw}\mathrm{err}_{vw}\leq\frac{1}{1-3\varepsilon_{1}}\cdot\frac{2\varepsilon_{1}}{y_{\emptyset}}|E_{\mathrm{adm}}|$. This proves the lemma. ∎
4.4 Construction of Solution to the Cluster LP Using Independently Sampled Clusters

### v3

@@ 3 Preclustering
Given a Correlation Clustering instance $(V,E^{+}\uplus E^{-})$, a preclustered instance is defined by a pair $({\mathcal{K}},E_{\mathrm{adm}})$, where $\mathcal{K}$} is a partition of $V$ (which can also be viewed as a clustering), and $E_{\mathrm{adm}}\subseteq\binom{V}{2}$ is a set of pairs such that for every $v\in E_{\mathrm{adm}}, uu$ and $v$ are not in a same set in $\mathcal{K}$}.
@@
We show that we can efficiently produce a preclustered instance for which there is a $(1+\epsilon)$-approximate good clustering. The rest of this section is dedicated to the proof of the following theorem:
@@
[…]
3.1 Constructing Atoms $\mathcal{K}$}
Our $\mathcal{K}$} is obtained from $\mathcal{C}$} by marking some vertices and creating singletons for them. We view $\mathcal{K}$} both as a clustering and as the set of atoms. The algorithm is described in Algorithm 5, where $\beta=0.1$.
Algorithm 5 Construction of $\mathcal{K}$}
1:for every non-singleton $C\in{\mathcal{C}}$ do
5:return $\mathcal{K}$}
Lemma 14.
Proof.
If a cluster $C$ has fewer than $\frac{\beta|C|}{3}$ marked vertices before Step 3, then the cost incurred by separating the marked vertices in $C$ is at most $\frac{2}{\beta}=O(1)$ times the cost of edges incident to these vertices in $\mathcal{C}$}. On the other hand, if $C$ has at least $\frac{\beta|C|}{3}$ marked vertices before Step 3, the cost of edges incident to $C$ in $\mathcal{C}$} is at least $\frac{\beta}{2}\cdot|C|\cdot\frac{\beta|C|}{3}\cdot\frac{1}{2}=\Omega(|C|^{2})$. The cost incurred by breaking $C$ into singletons is at most $|C|\choose 2$. As every edge is charged at most twice, the cost incurred by creating singletons for all marked vertices is at most $O(1)\cdot\mathrm{obj}({\mathcal{C}})$. So $\mathrm{obj}({\mathcal{K}})\leq O(1)\cdot{\mathrm{opt}}$ as $\mathrm{obj}({\mathcal{C}})\leq 3\cdot{\mathrm{opt}}$. ∎
Lemma 15.
For every non-singleton atom $K\in{\mathcal{K}}$, and every $u\in K$, we have $|N^{+}_{u}\triangle K|<\beta|K|$.
Proof.
Lemma 16.
Proof.
With Lemma 16, we now restrict ourselves to clusterings that do not break atoms. As a result, we can then assume all the edges between two vertices in a same atom $K\in{\mathcal{K}}$ are ++edges. The assumption only decreases $\mathrm{opt}$} and can only make the two properties of Theorem 13 harder to satisfy. We use $K_{u}$ for every $u\in V$ to denote the atom that contains $u$. We let $k_{u}=|K_{u}|$.
It is convenient for us to distribute the ++edges incident to an atom $K\in{\mathcal{K}}$ equally to the vertices in $K$. For every $u,v\in V$, we define
to be the probability that an edge between a random vertex in $K_{u}$ and a random vertex in $K_{v}$ is a ++edge. So $w_{uv}\in[0,1]$ for every pair $uv$; in particular, if $v\in K_{u}$, we have $w_{uv}=1$. We shall call $w_{uv}$ the weight of the edge $uv$. For any set $u\in V,V^{\prime}\subseteq V$, we define $w(u,V^{\prime}):=\sum_{v\in V^{\prime}}w_{uv}$ be the total weight of edges between $u$ and $V^{\prime}$. Let $w_{u}:=w(u,V)$ be the total weight of all edges incident to $u$. In this new instance, $w_{uv}$ fraction of edge $uv$ has a ++ sign, and the remaining $1-w_{uv}$ fraction has a - sign; $w_{u}$ is the total fractional number of ++edges incident to $u$, which can be treated as the ++degree of $u$. This new instance is equivalent to the original one. Till the end of this section we focus on this instance and call it the averaged instance to distinguish it from the original one.
3.2 Defining Admissible Edges $E_{\mathrm{adm}}$
In this section, we define the set $E_{\mathrm{adm}}$ of admissible edges, such that the following holds. There is a $(1+O(\varepsilon))$-approximate clustering ${\mathcal{C}}^{*}_{1}$ that does not break atoms, and all edges between two different clusters in ${\mathcal{C}}^{*}_{1}$ are admissible. We shall bound $|E_{\mathrm{adm}}|$ in terms of $\mathrm{obj}({\mathcal{K}})$, which is at most $O(1)\cdot{\mathrm{opt}}$. Notice that the construction of $E_{\mathrm{adm}}$ is a part of our algorithm for the proof of Theorem 13, but the construction of ${\mathcal{C}}^{*}_{1}$ is not. So, the latter procedure can depend on the optimum clustering $\mathcal{C}$$^{*}$.
We give the definition of $E_{\mathrm{adm}}$ upfront:
$E^{1}$ $:=\Big\{uv:\varepsilon w_{v}<w_{u}<\frac{w_{v}}{\varepsilon}\Big\},$ (6)
$E^{2}$ $:=\Big\{uv:K_{u}=K_{v}\text{ or }uv\in E^{1}\text{ and }\sum_{p\in N^{1}_{u}\cap N^{1}_{v}}w_{up}w_{vp}>\varepsilon(w_{u}+w_{v})\Big\},$ (7)
$E_{\mathrm{adm}}$ $:=\Big\{uv\in E^{2}:K_{u}\neq K_{v}\Big\}.$ (8)
Above, $uv$ is an unordered pair in $V$, with $u=v$ allowed. As a sanity check, notice that the conditions for the three definitions are symmetric between $u$ and $v$. Notice that $E_{\mathrm{adm}}\subseteq E^{2}\subseteq E^{1}$. We use $N^{1}_{u},N^{2}_{u}$ to denote the neighbor sets of $u$ in graphs $(V,E^{1})$ and $(V,E^{2})$ respectively. We use $d^{2}_{u}=|N^{2}_{u}|$ to denote the degree of $u$ in $(V,E^{2})$.
We then proceed to show the existence of a clustering ${\mathcal{C}}^{*}_{1}$ with desired properties.
Lemma 17.
Proof.
The inequality is by the condition of while loop.
Again, the second inequality is by the condition of the while loop.
We let ${\mathcal{C}}^{*}_{1}$ be the $(1+O(\varepsilon))$-approximate clustering satisfying the properties of Lemma 17.
Lemma 18.
Proof.
By Lemma 18 and the definition of $E^{1}$ in 6, we have
Corollary 19.
Lemma 20.
Proof.
The second inequality is by that $w_{vp},w_{up}\in[0,1]$. ∎
Corollary 19, Lemma 20 and the definition of $E^{2}$ in (7) imply
Corollary 21.
So, by the definition of $E_{\mathrm{adm}}$ in (8), ${\mathcal{C}}^{*}_{1}$ is a good clustering w.r.t the preclustered instance $({\mathcal{K}},E_{\mathrm{adm}})$.
We then proceed to bound $|E_{\mathrm{adm}}|$.
Lemma 22.
$d^{2}_{u}-k_{u}\leq O(\frac{1}{\varepsilon^{2}})(w_{u}-k_{u})$.
Proof.
The total number of admissible edges is
4 Solving Cluster LP Approximately
@@
We apply Theorem 13 to obtain a preclustered instance $({\mathcal{K}},E_{\mathrm{adm}})$, with the unknown good clustering ${\mathcal{C}}^{*}_{1}$. We can assume in the preclustered instance $({\mathcal{K}},E_{\mathrm{adm}})$, the edges between two different atoms $K$ and $K^{\prime}$ are all admissible, or all non-admissible. If one edge between them is non-admissible, we can change all other edges to non-admissible edges. This will not change the set of good clusterings, and it will decrease $|E_{\mathrm{adm}}|$.
@@
We define $K_{u}$ to be the atom that contains $u$, and $k_{u}=|K_{u}|$. We shall use $N_{\mathrm{adm}}(u)$ to be the set of vertices $v$ such that $uv\in E_{\mathrm{adm}}$; so $N_{\mathrm{adm}}(u)=N_{\mathrm{adm}}(v)$ if $v\in K_{u}$. We further process the good clustering ${\mathcal{C}}^{*}_{1}$ using the following procedure. It is not a part of our algorithm; it is only for analysis purpose.
@@ 4.1 Bounded Sub-Cluster LP Relaxation for Preclustered Instances
We form an LP relaxation aiming at finding the good clustering ${\mathcal{C}}^{*}_{1}$. In the LP, we have a variable $y^{s}_{S}$, for every $s\in[n]$, and $S\subseteq V$ of size at most $r$ (recall that $r=\Theta(1/\varepsilon^{12})$), that denotes the number of clusters in ${\mathcal{C}}^{*}_{1}$ of size $s$ containing $S$ as a subset. When $S\neq\emptyset$, there is at most one such cluster and thus $y^{s}_{S}\in\{0,1\}$ indicates if $S$ is a subset of a cluster of size $s$ in ${\mathcal{C}}^{*}_{1}$. For every $S\subseteq V$ of size at most $r$, let $y_{S}:=\sum_{s}y^{s}_{S}$ denote the number of clusters (of any size) in ${\mathcal{C}}^{*}_{1}$ containing $S$ as a subset. Again, if $S\neq\emptyset$, then $y_{S}\in\{0,1\}$ indicates if $S$ is a subset of a cluster in ${\mathcal{C}}^{*}_{1}$. For every $uv\in{V\choose 2}$, we have a variable $x_{uv}$ indicating if $u$ and $v$ are separated or not in ${\mathcal{C}}^{*}_{1}$. We call the LP the bounded sub-cluster LP relaxation, as we have variables indicating if a small set $S$ is a subset of a cluster or not.
@@
(9) gives the definition of $y_{S}$, (10) requires $u$ to be contained in some cluster, and (11) gives the definition of $x_{uv}$. (12) says if $y^{s}_{S}=1$, then there are exactly $s$ elements $u\in V$ with $y^{s}_{Su}=1$. (An exception is when $S=\emptyset$; the equality also holds in this case.) (13) is the non-negativity constraint. (14) and (15) follows from that ${\mathcal{C}}^{*}_{1}$ is a good clustering, and (16) follows from (A1). The left side of (17) is the number of clusters of size $s$ containing $S$ but does not contain any vertex in $T$ by inclusion-exclusion principle. So the inequality holds. This corresponds to a Sherali-Adams relaxation needed for the correlated rounding [RT12], see Lemma 24. The running time for solving the LP is $n^{O(r)}=n^{O(1/\varepsilon^{12})}$.
@@ 4.2 Sampling One Cluster Using LP Solution to the Bounded Sub-Cluster LP
We solve the bounded sub-cluster LP to obtain the $y$ and $x$ vectors. Given $y$, we can use the procedure construct-cluster described in Algorithm 6, to produce a random cluster $C$. Notice that steps 1 and 2 of the algorithm are well-defined due to equalities (9) and (12).
Algorithm 6 Construct-Cluster(y)(y)
@@
2:randomly choose a vertex $u\in V$, so that $u$ is chosen with probability $\frac{y^{s}_{u}}{sy^{s}_{\emptyset}}$; call $u$ the pivot
3:define a vector $y^{\prime}$ such that $y^{\prime}_{S}=\frac{y^{s}_{Su}}{y^{s}_{u}}$ for every $S\subseteq V$ of size at most r1r-1
@@
We define $\mathrm{err}^{s}_{vw|u}$ to be the error generated by the procedure when we choose $s$ as the cardinality and $u$ as the pivot:
[…]
@@
We prove the following two lemmas.
Lemma 25.
@@
Proof.
The probability is
$\sum_{s}\frac{y^{s}_{\emptyset}}{y_{\emptyset}}\sum_{u\in V^{\prime}}\frac{y^{s}_{u}}{sy^{s}_{\emptyset}}\cdot\frac{y^{s}_{uv}}{y^{s}_{u}}=\frac{1}{y_{\emptyset}}\sum_{s}\frac{1}{s}\sum_{u\in V^{\prime}}y^{s}_{uv}=\frac{1}{y_{\emptyset}}\sum_{s}y^{s}_{v}=\frac{1}{y_{\emptyset}}y_{v}=\frac{1}{y_{\emptyset}}.$
The second equality is by (12). The third and the last inequalities are by (9) and (10) respectively. ∎
Lemma 26.
@@
    $\Pr\left[\{v,w\}\subseteq C\right]\leq\frac{1}{y_{\emptyset}}\cdot y_{vw}+\mathrm{err}_{vw}$.
Proof.
We focus on the first statement. The probability that $v\in C$ and $w\notin C$ conditioned on $s$ is at most
$\sum_{u\in V}\frac{y^{s}_{u}}{sy^{s}_{\emptyset}}\cdot\left(\frac{1}{y^{s}_{u}}\cdot\big(y^{s}_{uv}-y^{s}_{uvw}\big)+\mathrm{err}^{s}_{vw|u}\right)$ $=\sum_{u\in V}\left(\frac{1}{sy^{s}_{\emptyset}}\cdot(y^{s}_{uv}-y^{s}_{uvw})+\frac{y^{s}_{u}}{sy^{s}_{\emptyset}}\cdot\mathrm{err}^{s}_{vw|u}\right)$
$=\frac{1}{y^{s}_{\emptyset}}(y^{s}_{v}-y^{s}_{vw})+\mathrm{err}^{s}_{vw}.$
To see the second equality, we apply (12) with $S=\{v\},\{w\}$ and $\{v,w\}$ respectively, and use the definition of $\mathrm{err}^{s}_{vw}$. Deconditioning on $s$, we have that the probability $vw$ is split by $C$ is at most
$\sum_{s}\frac{y^{s}_{\emptyset}}{y_{\emptyset}}\cdot\left(\frac{1}{y^{s}_{\emptyset}}(y^{s}_{v}-y^{s}_{vw})+\mathrm{err}^{s}_{vw}\right)$ $=\frac{1}{y_{\emptyset}}\sum_{s}(y^{s}_{v}-y^{s}_{vw})+\mathrm{err}_{vw}$
$=\frac{1}{y_{\emptyset}}(1-y_{vw})+\mathrm{err}_{vw}=\frac{1}{y_{\emptyset}}(x_{vw})+\mathrm{err}_{vw}.$
The second equality used that $\sum_{s}y^{s}_{v}=y_{v}=1$ and $\sum_{s}y^{s}_{vw}=y_{vw}=1-x_{vw}$.
For the second statement, the probability that $v,w\in C$ conditioned on $s$ is at most
$\sum_{u\in V}\frac{y^{s}_{u}}{sy^{s}_{\emptyset}}\cdot\left(\frac{1}{y^{s}_{u}}\cdot y^{s}_{uvw}+\mathrm{err}^{2}_{vw|u}\right)$ $=\sum_{u\in V}\left(\frac{1}{sy^{s}_{\emptyset}}\cdot y^{s}_{uvw}+\frac{y^{s}_{u}}{sy^{s}_{\emptyset}}\cdot\mathrm{err}^{s}_{vw|u}\right)$
$=\frac{1}{y^{s}_{\emptyset}}\cdot y^{s}_{vw}+\mathrm{err}^{s}_{vw}.$
Deconditioning on $s$, as before, results in the desired bound of $y_{vw}/y_{\emptyset}+\mathrm{err}_{vw}$. ∎
We can then bound the total unconditional error over all $vw$ pairs:
@@
Throughout the proof, we assume $u,v,w$ are all in $V$, vwvw and $uw$ are in $V\choose 2$.
Fix some $s\in[n],u\in V$ with $y^{s}_{u}>0$, and we now bound $\sum_{vw}\mathrm{err}^{s}_{vw|u}$. If $s=k_{u}$, then $C=K_{u}$; no errors will be created and the quantity is 0. Assume $s>k_{u}$. By (16), we have that $s>k_{u}+\varepsilon_{1}\cdot|N_{\mathrm{adm}}(u)|$, since otherwise we shall have $y^{s}_{u}=0$. By the second property of Lemma 24, we have $\sum_{vw}\mathrm{err}^{s}_{vw|u}\leq\frac{\varepsilon_{\mathrm{rt}}}{2}|N_{\mathrm{adm}}(u)|^{2}$. (Notice that if one of $v$ and $w$ is not in $N_{\mathrm{adm}}(u)$, then $\mathrm{err}^{s}_{vw|u}=0$.) Recall that $\varepsilon_{\mathrm{rt}}=\varepsilon_{1}^{2}$. Therefore,
@@
$=\frac{\varepsilon_{1}}{2}\cdot|N_{\mathrm{adm}}(u)|\cdot\sum_{v\in N_{\mathrm{adm}}(u)}\frac{y^{s}_{uv}}{y^{s}_{u}}=\frac{\varepsilon_{1}}{2}\cdot\sum_{v,w\in N_{\mathrm{adm}}(u)}\frac{y^{s}_{uv}}{y^{s}_{u}}.$ (18)
The first equality is by (12) and $y^{s}_{uv}=y^{s}_{u}$ for every $v\in K_{u}$. (To see this, notice that $y^{s}_{uv}\leq y^{s}_{u}$ is implied by (17). We have $y_{uv}=\sum_{s}y^{s}_{uv}$, $y_{u}=\sum_{s}y^{s}_{u}$, and $y_{uv}=y_{u}=1$ if $v\in K_{u}$.) The second equality is obtained by replacing $|N_{\mathrm{adm}}(u)|$ with $\sum_{w\in N_{\mathrm{adm}}(u)}1$.
@@
$=\varepsilon_{1}\sum_{uw\in E_{\mathrm{adm}}}\Pr[C\cap\{u,w\}\neq\emptyset\mid s].$ (19)
The first inequality is by (18). For the second inequality in the above sequence, notice that we view $uw$ as an unordered pair. Thus we keep both $y^{s}_{uv}$ and $y^{s}_{vw}$ in the numerator. To see the third inequality, notice that $\frac{y^{s}_{uv}}{y^{s}_{v}}=\Pr[u\in C|s,v\text{ is pivot}]\leq\Pr[C\cap\{u,w\}\neq\emptyset|s,v\text{ is pivot}]$. The same inequality holds for $\frac{y^{s}_{vw}}{y^{s}_{v}}$.
@@
The first inequality is by (19). To see the second inequality, we notice that $C\cap\{u,w\}\neq\emptyset$ is the union of the 3 disjoint events: $u\in C$ and $w\notin C$, $u\notin C$ and $w\in C$, and $\{u,w\}\subseteq C$. By Lemma 26, we have $\Pr[C\cap\{u,w\}\neq\emptyset]\leq\frac{2x_{uw}+y_{uw}}{y_{\emptyset}}+3\cdot\mathrm{err}_{uw}\leq\frac{2}{y_{\emptyset}}+3\cdot\mathrm{err}_{uw}$. So, we have $\sum_{vw}\mathrm{err}_{vw}\leq\frac{1}{1-3\varepsilon_{1}}\cdot\frac{2\varepsilon_{1}}{y_{\emptyset}}|E_{\mathrm{adm}}|$. This proves the lemma. ∎
4.3 Using Independently Sampled Clusters to Construct Cluster LP Solution